OpenStudy (anonymous):

Write a justification for each step AB=BC 5y+6=2y+21 3y+6=21 3y=15 y=5 *I have to reason with it*

OpenStudy (anonymous):

Do you know what substitution is

OpenStudy (anonymous):

???????????????/

OpenStudy (anonymous):

I edit the equation

OpenStudy (anonymous):

for the problem: 5y+6=2y+21 since y=5 you do 5(5)+6=2(5)+21 25 +6=10+21

OpenStudy (anonymous):

for the problem: 3y+6=21 do the same